OpenStudy (anonymous):

The stopping distance d of a car after the brakes are applied varies directly as the square of the speed r. If a car traveling 50 mph can stop in 140ft,. How many feet will it take the same car to stop when it is traveling 40 mph?

OpenStudy (anonymous):

140=k*50^2 140=k*2500 k=140/2500 = 0.056 d=.056r^2 d=.056*40^2 d=89.6ft